J.D. Power and Associates says that 60 percent of car buyers now use the Internet for research and price comparisons. (a) Find the probability that in a sample of 8 car buyers, all 8 will use the Internet; (b) at least 5; (c) more than 4. (d) Find the mean and standard deviation of the probability distribution. (e) Sketch the PDF (using Excel or Appendix A) and describe its appearance (e.g., skewness). (Data are from J. Paul Peter and Jerry C. Olson, Consumer Behavior and Marketing Strategy, 7th ed. [McGraw-Hill/Irwin, 2005], p. 188.)
Answer by stanbon(75887)

J.D. Power and Associates says that 60 percent of car buyers now use the Internet for research and price comparisons.
(a) Find the probability that in a sample of 8 car buyers, all 8 will use the Internet;
Ans: 0.6^8
---------------
(b) at least 5
P(5<= x <= 8) = 1 - binomcdf(8,0.6,4) = 0.5941
----------------------------
(c) more than 4.
P(5<= x <=8) = 0.5941
----------------------------
(d) Find the mean and standard deviation of the probability distribution.
mean = np = 8*0.6 = 4.8
std = sqrt(npq) = sqrt(4.8*0.4) = 1.3856....
------------------------------
(e) Sketch the PDF (using Excel or Appendix A) and describe its appearance (e.g., skewness).
Flater to the left of the mean because p > 0.5
Other?
==========================================
Cheers,
Stan H.
==============================